The florist advertises roses for $2 a bud and carnations for $0.75 a flower. If Kathleen pays $20.50,excluding tax, for a bouquet of fourteen flowers, how many of each flower are in the bouquet?
Answer by fractalier(6550)

Call x the number of roses...so that the number of carnations is 14-x. Then we write
2x + .75(14-x) = 20.50
2x + 10.50 - .75x = 20.50
1.25x = 10
x = 8 roses
14-x = 6 carnations